# On the Effects of Modeling Errors on Distributed Continuous-time Filtering

### v2

@@ Abstract
In practical scenarios, obtaining precise models is rarely feasible, and modeling errors are widespread. The modeling errors involve deviations in the state matrix, deviations in the measurement matrix, and mismatched noise covariances. Such modeling errors have the potential to result in a decline in the filter performance and even lead to system failure. Several studies have explored the performance of continuous-time filters for a single sensor when confronted with modeling errors. T. Nishimura [12] established a conservative design criterion to ensure the upper bound of the estimation error covariance for continuous-time systems with incorrect noise covariances. Fitzgerald [4] investigated into the divergence of the Kalman filter, highlighting the mean square error may become unbounded due to the incorrect process noise covariance. Toda and Patel [19] derived performance index bounds and mean square error bounds for the suboptimal filters, considering the presence of modeling errors. Sangsuk-Iam and Bullock [15] analyzed the behavior of the continuous-time Kalman filter under mismatched noise covariances. Additionally, some existing literature on state estimation addresses arbitrary noise scenarios [17, 5, 9]. However, these studies primarily focus on the single-sensor systems. When shifting the focus to the distributed filters, numerous challenges arise. These challenges include dealing with the new distributed filter structure, addressing the coupling terms, and handling the behavior of the consensus parameter. This paper will investigate the effect of the modeling errors on the distributed continuous-time filters.
@@
    We define two performance indices, namely the nominal performance index and the estimation error covariance, as metrics to assess the performance of the filter. By utilizing the nominal performance index and the norm of the modeling deviations, the bounds of the estimation error covariance are determined (Theorem 1). In particular, the asymptotic decay behavior of this bound in terms of $\gamma_{u}$ is revealed. These results play an important role in evaluating the the estimation error covariance by utilizing the nominal models in the presence of the modeling errors.
@@ 1 Problem Statement
where $t$ is the time index, $x(t)\in\mathcal{R}^{n}$ is the system state, $y_{i}(t)\in\mathcal{R}^{m_{i}}$ is the measurement of sensor $i$, $A\in\mathcal{R}^{n\times n}$ is the state matrix, $C_{i}\in\mathcal{R}^{m_{i}\times n}$ is the measurement matrix of sensor $i$, $\omega(t)\in\mathcal{R}^{n}$ is the process noise, and $\nu_{i}(t)\in\mathcal{R}^{m_{i}}$ is the measurement noise of sensor $i$. It is assumed that $\omega(t)$ and $\nu_{i}(t)$ are zero-mean white noise processes, and these noise processes are uncorrelated with each other and with the initial state $x(0)$, i.e., $\mathbb{E}\{\omega(t)\omega^{T}(\tau)\}=Q\delta(t-\tau)$, $\mathbb{E}\{\nu_{i}(t)\nu^{T}_{i}(\tau)\}=R_{i}\delta(t-\tau),~{}i=1,\ldots,N,$ $\mathbb{E}\{\omega(t)\nu^{T}_{i}(\tau)\}=0,~{}i=1,\ldots,N,$ and $\mathbb{E}\{\nu_{i}(t)\nu^{T}_{j}(\tau)\}=0,~{}i,j=1,\ldots,N,i\neq j$, $Q\in\mathcal{R}^{n\times n}$ and $R_{i}\in\mathcal{R}^{m_{i}\times m_{i}}$ denote the process noise covariance and the measurement noise covariance of sensor $i$, respectively, and $\delta(t)$ represents the Dirac delta function, given by $\delta(t-\tau)=1$ if $t=\tau$ and 0 otherwise. The communication topology of sensor networks is described by the graph theory in Appendix A.
@@
where $K_{i}=NP(\infty)C^{T}_{i}R^{-1}_{i}$, $P(\infty)$ is the solution of
@@ 2.1 Nominal Distributed Kalman Filter
where $K_{i,u}=NP_{u}(\infty)C^{T}_{i,u}R^{-1}_{i,u}$, $P_{u}(\infty)$ is the solution of
@@ 3.1 Bounds of the Estimation Error Covariance
For Assumption 3, if $F_{d,u}\neq 0$ and $A$ is unstable, the estimation error covariance will diverge, making it meaningless to investigate the corresponding result. Hence, Assumption 3 has been made to provide a unified expression and analysis.
@@
For the distributed filter with the nominal model, it is desired that the steady-state estimation error covariance $\bar{\Sigma}_{e}$ can be evaluated by the nominal performance index $\bar{\Sigma}_{u}$. Consequently, the bounds of the steady-state estimation error covariance are given in the following. Since $\mathcal{A}_{u}$ incorporates $\gamma_{u}$, and $\gamma_{u}$ may tend to infinity. Hence, it is crucial to identify the effect of $\gamma_{u}$.
@@ 3.2 Divergence Analysis
When the communication topology is mismatched, it can lead to the divergence of the estimation error covariance. This may violate the condition presented in Proposition 2, resulting in the instability of $\mathcal{A}_{u}$. When the mismatched topology is denoted as $\mathcal{L}_{u}$, we can use the lower bound of $\lambda_{N-1}(\mathcal{L}_{u})$ to compute $\gamma_{u_{0}}$ in (19) to guarantee the filter performance.
@@ 3.3 Relations between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$
Consider the system (1) and the nominal distributed filter (5). Suppose Assumptions 1-2 hold. If $\gamma_{u}\geq\gamma_{u_{0}}$ (as defined in (20)), $\Delta A=0$, and $\Delta C_{i,u}=0,i=1,\ldots,N$, it holds
@@ 4 Simulations
In this section, numerical experiments are conducted on vehicle tracking to validate the effectiveness of the theoretical results. The sensor network used in the experiments consists of six sensors, and its communication topology is illustrated in Fig. 1.
@@
Consider that the vehicle is moving in a plane. Let $x(t)=[x^{T}_{1}(t),x^{T}_{2}(t),x^{T}_{3}(t),x^{T}_{4}(t)]^{T}$ denote the state, where $x_{1}(t)$ and $x_{3}(t)$ are the horizontal velocity and the vertical velocity, respectively, and $x_{2}(t)$ and $x_{4}(t)$ are the horizontal position and the vertical position, respectively. The vehicle dynamics is described by $A=[\bar{A}_{11},0;0,\bar{A}_{22}]$, where $\bar{A}_{11}=\bar{A}_{22}=[0,0;1,0]$, The process noise covariance is given as   $Q=\text{diag}\{0.03,0.03,0.03,0.03\}$. Two kinds of sensors are deployed in the sensor network, and their measurement matrices are defined as $C_{1}=[0,1,0,0]$, $C_{2}=[0,2,0,0]$, $C_{3}=[0,3,0,0]$, $C_{4}=[0,0,0,1]$, $C_{5}=[0,0,0,2]$, and $C_{6}=[0,0,0,3]$. The measurement noise covariances are set as $R_{i}=0.2,~{}i=1,\ldots,6$. The initial state is $x(0)=[0.2;1;0.2;1]$, and the initial covariance is $\Sigma_{i,e}(0)=\text{diag}\{0.1,0.1,0.1,0.1\}$. The Monte Carlo method is adopted to evaluate the performance of the distributed filter, and the mean square error (MSE) is utilized as a metric, given by $\text{MSE}(t)=\frac{1}{MN}\sum^{N}_{i=1}\sum^{M}_{l=1}\|\hat{x}^{(l)}_{i,u}(t)-x^{(l)}(t)\|^{2}_{2},$ where $N$ is the sensor number and $M$ is the trail number.
@@
    As the consensus parameter $\gamma_{u}$ increases, five indices decline. Moreover, both MSE and $\text{Tr}(\bar{\Sigma}_{e})$ are constrained by the upper bound 1 and the upper bound 2, and the asymptotic decay of the upper bound is demonstrated, as established Theorem 1.

### v4

@@ Abstract
In practical scenarios, obtaining precise models is rarely feasible, and modeling errors are widespread. The modeling errors involve deviations in the state matrix, deviations in the measurement matrix, and mismatched noise covariances. Such modeling errors have the potential to result in a decline in the filter performance and even lead to system failure. Several studies have explored the performance of continuous-time filters for a single sensor when confronted with modeling errors. T. Nishimura [12] established a conservative design criterion to ensure the upper bound of the estimation error covariance for continuous-time systems with incorrect noise covariances. Griffin and Sage [6] conducted a thorough sensitivity analysis of filtering and smoothing algorithms, exploring both large and small-scale scenarios, in the presence of the modeling errors. Fitzgerald [4] investigated the divergence of the Kalman filter, highlighting the mean square error may become unbounded due to the incorrect process noise covariance. Toda and Patel [19] derived performance index bounds and mean square error bounds for the suboptimal filters, considering the presence of modeling errors. Sangsuk-Iam and Bullock [15] analyzed the behavior of the continuous-time Kalman filter under mismatched noise covariances. Additionally, some existing literature on state estimation addresses arbitrary noise scenarios [17, 5, 9]. However, these studies primarily focus on the single-sensor systems. When shifting the focus to the distributed filters, numerous challenges arise. These challenges include dealing with the new distributed filter structure, addressing the coupling terms, and handling the behavior of the consensus parameter. This paper will investigate the effect of the modeling errors on the distributed continuous-time filters.
@@
    We define two performance indices, namely the nominal performance index and the estimation error covariance, as metrics to assess the performance of the filter. By utilizing the nominal performance index and the norm of the modeling deviations, the bounds of the estimation error covariance are determined (Theorem 1). In particular, the asymptotic decay behavior of the bounds in terms of the nominal consensus parameter $\gamma_{u}$ is revealed. These results play an important role in evaluating the estimation error covariance by utilizing the nominal models in the presence of the modeling errors.
@@ 1 Problem Statement
where $t$ is the time index, $x(t)\in\mathcal{R}^{n}$ is the system state, $y_{i}(t)\in\mathcal{R}^{m_{i}}$ is the measurement of sensor $i$, $A\in\mathcal{R}^{n\times n}$ is the state matrix, $C_{i}\in\mathcal{R}^{m_{i}\times n}$ is the measurement matrix of sensor $i$, $\omega(t)\in\mathcal{R}^{n}$ is the process noise, and $\nu_{i}(t)\in\mathcal{R}^{m_{i}}$ is the measurement noise of sensor $i$. It is assumed that $\omega(t)$ and $\nu_{i}(t)$ are zero-mean white noise processes, and these noise processes are uncorrelated with each other and with the initial state $x(0)$, i.e., $\mathbb{E}\{\omega(t)\omega^{T}(\tau)\}=Q\delta(t-\tau)$, $\mathbb{E}\{\nu_{i}(t)\nu^{T}_{i}(\tau)\}=R_{i}\delta(t-\tau),~{}i=1,\ldots,N,$ $\mathbb{E}\{\omega(t)\nu^{T}_{i}(\tau)\}=0,~{}i=1,\ldots,N,$ and $\mathbb{E}\{\nu_{i}(t)\nu^{T}_{j}(\tau)\}=0,~{}i,j=1,\ldots,N,i\neq j$, $Q\in\mathcal{R}^{n\times n}$ and $R_{i}\in\mathcal{R}^{m_{i}\times m_{i}}$ denote the process noise covariance and the measurement noise covariance of sensor $i$, respectively, and $\delta(t)$ represents the Dirac delta function, given by $\delta(t-\tau)=1$ if $t=\tau$ and 0 otherwise. The communication topology of sensor networks is described by using the graph theory, as detailed in Appendix A.
@@
where $K_{i}=NP(\infty)C^{T}_{i}R^{-1}_{i}$, $\gamma$ is the consensus parameter, $P(\infty)$ is the solution of
@@ 2.1 Nominal Distributed Kalman Filter
where $K_{i,u}=NP_{u}(\infty)C^{T}_{i,u}R^{-1}_{i,u}$, $\gamma_{u}$ is the nominal consensus parameter, $P_{u}(\infty)$ is the solution of
@@ 3.1 Bounds of the Estimation Error Covariance
For Assumption 3, if $F_{d,u}\neq 0$ and $A$ is unstable, the estimation error covariance will diverge, making it meaningless to investigate the corresponding result. Assumption 3 is introduced to provide a unified framework for analysis, covering both cases where $F_{d,u}\neq 0$ and $F_{d,u}=0$. The case $F_{d,u}=0$ primarily reflects the scenarios where $A_{u}=A$ and $C_{i,u}=C_{i}$. For this case, the following results highlight the effect of mismatched noise covariances on the filter’s performance.
@@
For the distributed filter with the nominal model, it is desired that the steady-state estimation error covariance $\bar{\Sigma}_{e}$ can be evaluated by the nominal performance index $\bar{\Sigma}_{u}$. Consequently, the bounds of the steady-state estimation error covariance are given in the following. Since $\mathcal{A}_{u}$ incorporates $\gamma_{u}$, and $\gamma_{u}$ may tend to infinity. It is crucial to identify the effect of $\gamma_{u}$.
@@ 3.2 Divergence Analysis
When the communication topology is mismatched, it can lead to the divergence of the estimation error covariance. This may violate the condition presented in Proposition 2, resulting in the instability of $\mathcal{A}_{u}$. When the mismatched topology is denoted as $\mathcal{L}_{u}$, we can use the lower bound of $\lambda_{N-1}(\mathcal{L}_{u})$ to compute $\bar{\gamma}_{u_{0}}$ in (19), ensuring that $\mathcal{A}_{u}$ is Hurwitz stable by choosing $\gamma_{u}>\bar{\gamma}_{u_{0}}$.
@@ 3.3 Relations between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$
Consider the system (1) and the nominal distributed filter (5). Suppose Assumptions 1-2 hold. If $\gamma_{u}\geq\gamma_{u_{0}}$, $\Delta A=0$, and $\Delta C_{i,u}=0,i=1,\ldots,N$, it holds
@@ 4 Simulations
This section conducts numerical experiments on vehicle tracking to validate the effectiveness of the theoretical results regarding the effects of modeling errors on the filter performance. Consider the vehicle moving on a plane. In engineering applications, we primary focus on the vehicle’s positions and velocities. Let $x(t)=[x^{T}_{1}(t),x^{T}_{2}(t),x^{T}_{3}(t),x^{T}_{4}(t)]^{T}$ denote the state, where $x_{1}(t)$ and $x_{3}(t)$ represent the horizontal and vertical velocities, respectively, and $x_{2}(t)$ and $x_{4}(t)$ represent the horizontal and vertical positions, respectively. The vehicle dynamics is described by $A=[\bar{A}_{11},0_{2\times 2};0_{2\times 2},\bar{A}_{22}]$, where $\bar{A}_{11}=\bar{A}_{22}=[0,0;1,0]$. This represents the vehicle undergoing uniform liner motion with noise disturbances. The sensor network consists of six sensors to measure the vehicle’s state, and its communication topology is illustrated in Fig. 1.
@@
The process noise covariance is given as   $Q=\text{diag}\{0.03,0.03,0.03,0.03\}$. Two kinds of sensors are deployed in the sensor network, and their measurement matrices are defined as $C_{1}=[0,1,0,0]$, $C_{2}=[0,2,0,0]$, $C_{3}=[0,3,0,0]$, $C_{4}=[0,0,0,1]$, $C_{5}=[0,0,0,2]$, and $C_{6}=[0,0,0,3]$. The measurement noise covariances are set as $R_{i}=0.2,~{}i=1,\ldots,6$. The initial state is $x(0)=[0.2;1;0.2;1]$, and the initial covariance is $\Sigma_{i,e}(0)=\text{diag}\{0.1,0.1,0.1,0.1\}$. The Monte Carlo method is adopted to evaluate the performance of the distributed filter, and the mean square error (MSE) is utilized as a metric, given by $\text{MSE}(t)=\frac{1}{MN}\sum^{N}_{i=1}\sum^{M}_{l=1}\|\hat{x}^{(l)}_{i,u}(t)-x^{(l)}(t)\|^{2}_{2},$ where $N$ is the sensor number and $M$ is the trail number.
@@
    As the consensus parameter $\gamma_{u}$ increases, five indices decline. Moreover, both MSE and $\text{Tr}(\bar{\Sigma}_{e})$ are constrained by the upper bound 1 and the upper bound 2, and the asymptotic decay of the upper bound is demonstrated, as established in Theorem 1.